Estimate the ΔH_{reaction}​reaction​​ for the reaction shown below using data from table 9.3. Show calculations and answer should be in units of kJ/mol and reported to the ones place.
Answer #1

For the reaction

We have,

\DeltaHreaction = \sumBond energy of reactants-\sumBond energy of products

= [4(C-H) + (C=C) + (O = O)] - [ 4(C-H) + 2(C = O)] KJmol-1

= [ (4 x 412) + 612 + 497] - [ (4 x 412) + (2 x 743) ] kJmol-1

= [1648 + 612 + 497] - [1648 + 1486] kJmol-1

= 2757 - 3134= -377 kJmol-1